# SESR-Eval: Dataset for Evaluating LLMs in the Title-Abstract Screening of Systematic Reviews

### v1

@@ Abstract
Index Terms:
Title-abstract screening, Large language models, Systematic reviews, Dataset
@@ I Introduction
Systematic reviews (SRs) are a research method used to identify and interpret relevant research on a particular topic  [1]. This method is widely used in software engineering (SE) research. For example, more than a hundred SRs on software testing have been published to date [2]. One critical step in the SR process is the screening of primary studies, where each candidate study is rigorously evaluated against a set of predefined criteria [1]. The screening of primary studies in SRs typically involves two parts – (1) title-abstract screening and (2) full-text screening. Title-abstract screening allows for rapid filtering of relevant primary studies. Studies that are clearly irrelevant can be excluded, and the included studies can then be screened in-depth by full-text screening  [1]. This step is both time-consuming and prone to error, which has motivated recent research on the automation of the primary study screening process using large language models (LLMs) in SE  [3, 4] and more extensively in biomedical and medical domains [5, 6, 7, 8, 9, 10].
While prior studies show promising results for title-abstract screening in SE  [3, 4], they suffer from a small number of secondary studies and LLMs under evaluation. One study [3] evaluated performance of two LLMs in a single secondary study, while another study [4] evaluated performance of a single LLM with two secondary studies. Performance evaluations based on small datasets risk introducing bias, making it difficult to assess the utility of the solution.
@@ II-A Automating the Screening Process of Systematic Reviews
[3] SE 2 Binary 1 1,306
[4] SE 1 Likert 2 582
[5] Biomed 9 Binary 3 505
[6] Biomed 3 Likert 10 38,426
[7] Med 1 Binary 5 22,666
[8] Med 18 Binary 3 6,217
[9] Med 2 Binary 2 4,527
[10] Med 2 Binary 6 24,307
@@
Table I summarizes related work in this area. In the SE domain, two studies [3, 4] investigated the use of LLMs for screening with two different evaluation methods and 1,888 primary studies in total. In contrast, the biomedical and medical domains have seen more research on the topic [5, 6, 7, 8, 9, 10]. Studies outside of SE have used significantly larger sample of both secondary and primary studies.
@@ II-B Benchmarking Large Language Models
AGIEval [11, 12] Knowledge & reasoning
ARC-Challenge [13] Knowledge & reasoning
BIG-Bench-Hard [14] Knowledge & reasoning
GPQA [15], Diamond Knowledge & reasoning
HellaSwag [16] Knowledge & reasoning
MMLU [11, 17] Knowledge & reasoning
GSM8K [11, 18] Mathematical & logical reasoning
MATH [11, 19] Mathematical & logical reasoning
MathVista (testmini) [20] Mathematical & logical reasoning
MGSM [21] Mathematical & logical reasoning
MMMU [22] Mathematical & logical reasoning
HumanEval [11, 23] Programming & code generation
MBPP [24] Programming & code generation
WikiSQL [25] Programming & code generation
DROP [26] Multimodal & visual reasoning
AI2D, test [27] Multimodal & visual reasoning
ChartQA [28] Multimodal & visual reasoning
MMBench [29] Multimodal & visual reasoning
@@ III Methods
Fig. 1 shows an overview of how the dataset was created and how the experiments used to answer our research questions were conducted. We followed the SIGSOFT Empirical Standards for Software Engineering Research [30] as a framework for conducting this study. These standards provide guidance on reporting benchmarking studies.
@@ III-A Research questions
  • RQ4: What is the combined effect of LLMs and secondary studies on the screening results? Motivation: Understanding the combined effect of LLMs and secondary studies is important, if the certain LLMs perform significantly better in screening on specific secondary studies.
@@ III-B Creating the benchmark dataset
This subsection provides an answer for RQ1. Fig. 1 visualizes the benchmark dataset creation process. We used a prior systematic mapping [31] as our source for research artifacts, which lists 169 research artifacts.
@@
From the data extraction (Table III), the ”Screening results” information is used to include or exclude research artifacts. We double-coded 20 randomized research artifacts regarding whether the secondary study contains screening data and that data is suitable for our experiment. We had an agreement of 90% and 95%, corresponding to Krippendorff’s alphas [32] of 0.7860.7860.786 and 0.8310.8310.831 respectively. The rest of the artifacts were then investigated by the first author.
@@
After the criteria retrieval, we generated unique identifiers for each secondary study in the dataset to distinguish them when running benchmarks. Finally, we combined the primary studies from each secondary study to form the final SESR-Eval dataset. Papers with missing titles or abstracts were excluded from the final dataset, as title-abstract screening is meant to be conducted with non-empty titles and abstracts. The final dataset is available in the research artifact of this study [33].
@@ III-C1 Creating the zero-shot prompt template and constructing the primary study prompts
Before constructing the primary study prompts, we created the zero-shot prompt template ( 8), in Fig.  2. As a starting point for the prompt template, we used the zero-shot prompt from a SE paper, which studied title-abstract screening using LLMs [3]. Next, we added Likert scale decisions, used by Felizardo et al. [4], for each criterion (include or exclude) being met, and the relevance of the whole study:
@@
Finally, we added a floating point decision, which has a value between 0.0000.0000.000 and 1.0001.0001.000. A score of 1.0001.0001.000 means the paper (or criterion) is included with high probability; secondly - a score of 0.0000.0000.000 means the paper (or criterion) is most likely excluded. Our goal was to give the LLM as detailed prompt as possible, to reduce the chance of hallucinations and improve the overall task performance. We detailed each assessment type and emphasized that the LLM is required to produce all three types of assessments for the overall relevance of the study and for each individual inclusion-exclusion criteria.
@@ III-C2 Benchmarking title-abstract screening performance with LLMs
To study the screening performance of the LLMs with the larger dataset (RQ2) and across secondary studies (RQ3), we benchmarked the title-abstract screening performance with LLMs listed in Table IV. We aimed to include a diverse set of (a) reasoning, (b) non-reasoning, (c) commercial and (d) open-source LLMs, which publish their model weights or model source code online. For the reasoning models, we selected a commercial and an open-source LLM. Second, for non-reasoning models, we selected four commercial and two open-source LLMs. LLMs from these vendors are commonly used in SE research and are used in prior work [3, 4, 34, 35, 8]. Similarly, as the performance of neural networks with a relatively small number of parameters, sometimes also called small language models (SMLs) can achieve competitive performance [36], we decided to add two models that could be considered such to better understand time and token efficiency related to model size. In all of our benchmark experiments, we used the following system prompt:
@@
With the LLMs, we used structured JavaScript Object Notation (JSON) response format instead of the traditional textual response. Using a structured response format allows for a predictable output token count, programmatic integration and faster evaluation of the LLM output [37, 11]. The response format we used is described in Table  V and available in the research artifact [33].
@@
To compare the performance of title-abstract screening between LLMs and secondary studies (RQ4), we fit a logistic regression model predicting the correctness of screening decisions based on the LLM and the primary study. The logistic regression model was fitted and the odds ratios were calculated using the R programming language and the methods $glm$ and $coef$ included in the base R distribution version 4.3.34.3.34.3.3. For calculating Krippendorff’s Alpha, we used the R-package $irr$ [38] version 0.84.1. We also calculated the precision, recall and the F1 score for the tested LLMs and all primary studies and secondary studies from the benchmark set, using Python’s scikit-learn library [39]. The code for these is given in our research artifact.
For all LLMs, we set the temperature parameter 555 https://platform.openai.com/docs/api-reference/completions/create to zero and the top_p5 parameter to 0.10.10.1 – configuration used in related works [3, 4]. The temperature parameter controls how deterministic the LLM’s output is. The lower the temperature value, the more the LLM attempts to always output the same tokens for the same prompt, regardless of the number of calls to the LLM. Secondly, the top_p parameter controls the top n-tokens that are considered in the output. Setting top_p to 0.10.10.1 means that the LLM attempts to sample from the top 10% most probable tokens, which guides the LLM to output more relevant tokens.
@@ III-D How does the time, token efficiency, and cost vary across different LLMs?
This subsection provides background for the last research question (RQ5). To evaluate the time, token efficiency and cost of the LLMs, we analyzed the latency and input and output token counts for each of the LLMs using a custom Python script. The latency and token information is returned by the LLMs’ API. Latency was measured as the duration of sending a single request to the LLM and receiving its response. We used the 95th percentile [40] of response times to evaluate performance under typical conditions. The total token count of the LLM is a combination of the input (=prompt) and output tokens, where reasoning models are expected to output additional ”reasoning tokens”, which are aggregated to the output tokens. Providers, such as OpenAI, cache input tokens to reduce costs  [^6] https://platform.openai.com/docs/guides/prompt-caching. Costs were computed from the token count, multiplied by the LLMs price per 1M tokens.
@@ IV Results
[41] 1,194 742 451 62:38 0 Information and Software Technology 4 4 SE Process
[42] 223 191 32 86:14 1 Information and Software Technology 3 3 SE Process
[43] 458 146 312 32:68 53 Information and Software Technology 3 4 Software Testing
[44] 322 171 151 53:47 0 Software and Systems Modeling 3 6 SE Models and Methods
[45] 10,454 547 9,907 5:95 759 Journal of Systems and Software 4 6 Software Maintenance
[46] 4,671 54 4,617 1:99 968 Empirical Software Engineering 1 1 Software Design
[47] 1,733 169 1,564 10:90 82 Information and Software Technology 13 0 Software Testing
[48] 606 144 462 24:76 47 Information and Software Technology 3 3 Software Maintenance / Operations
[49] 2,541 74 2,467 3:97 281 Software and Systems Modeling 6 4 Software Architecture
[50] 540 36 504 7:93 219 Journal of Systems and Software 3 6 Software Requirements
[51] 1,741 290 1,451 17:83 717 Automated Software Engineering 2 6 Software Quality
[52] 2,353 73 2,280 3:97 274 Journal of Systems and Software 2 7 Software Quality
[53] 69 69 0 100:0 0 Journal of Systems and Software 3 4 Software Configuration Management
[54] 3,194 234 2,960 7:93 446 Empirical Software Engineering 7 7 SE Professional Practice
[55] 327 153 174 47:53 1 Journal of Systems and Software 5 6 SE Models and Methods
[56] 113 55 58 49:51 2 Information and Software Technology 3 3 Software Quality
[57] 731 522 209 72:28 236 Information and Software Technology 3 7 Software Testing
[58] 167 56 111 34:66 0 Information and Software Technology 6 8 SE Models and Methods
[59] 318 44 274 14:86 17 Journal of Systems and Software 4 7 Software Maintenance / SE Economics
[60] 5,454 223 5,231 4:96 1,268 Information and Software Technology 2 3 SE Management / SE Professional Practice
[61] 1,512 161 1,351 11:89 25 IEEE Transactions on Software Engineering 4 0 SE Models & Methods
[62] 318 187 131 59:41 0 Journal of Systems and Software 1 6 SE Models & Methods
[63] 802 127 675 16:84 21 ACM Trans. Softw. Eng 5 5 SE Professional Practice
[64] 105 105 0 100:0 1 Information and Software Technology 3 4 Software Testing
@@
Yes, 24 secondary studies were found with title-abstract screening data. The process of how the dataset was created is explained in Section III-B. In total, we reviewed 169 research artifacts, from which 11 (6.5%) were inaccessible due to reasons such as dead links, empty repositories or pages requiring authorization. Ultimately, 24 of the 169 artifacts (14.2%) were suitable for our benchmark dataset. The secondary studies selected for the benchmark dataset are shown in Table VI. The number of primary studies per secondary study varies significantly in the dataset — ranging from under a hundred papers to over 10,000 papers. Similar variations exist in the included/excluded ratio of papers. Two secondary studies contained only included studies, resulting in an I/E ratio of 100:0100:0100 : 0, while at the other end of the spectrum, one study included only about one percent of papers, with an I/E ratio of 1:991:991 : 99. On average, looking at the ratio of included and excluded studies in the dataset, 12% of primary studies were labeled as included and 88% as excluded. The median number of the inclusion and exclusion criteria across secondary studies was 333 and 4.54.54.5, respectively. Based on the Software Engineering Body of Knowledge (SWEBOK) [65], the dataset covers a wide spectrum of SE domains.

### v2

@@ Abstract
publicationid: pubid: 979-8-3315-9147-2/25/\$31.00 ©2025 IEEE
@@ I Introduction
Systematic reviews (SRs) are a research method used to identify and interpret relevant research on a particular topic  [kitchenhamGuidelinesPerformingSystematic2007]. This method is widely used in software engineering (SE) research. For example, more than a hundred SRs on software testing have been published to date [garousiSystematicLiteratureReview2016]. One critical step in the SR process is the screening of primary studies, where each candidate study is rigorously evaluated against a set of predefined criteria [kitchenhamGuidelinesPerformingSystematic2007]. The screening of primary studies in SRs typically involves two parts – (1) title-abstract screening and (2) full-text screening. Title-abstract screening allows for rapid filtering of relevant primary studies. Studies that are clearly irrelevant can be excluded, and the included studies can then be screened in-depth by full-text screening  [kitchenhamGuidelinesPerformingSystematic2007]. This step is both time-consuming and prone to error, which has motivated recent research on the automation of the primary study screening process using large language models (LLMs) in SE  [huotalaPromiseChallengesUsing2024a, felizardoChatGPTApplicationSystematic2024] and more extensively in biomedical and medical domains [liEvaluatingEffectivenessLarge2024, dennstadtTitleAbstractScreening2024, tranSensitivitySpecificityAvoidable2023, delgado-chavesTransformingLiteratureScreening2025, matsuiHumanComparableSensitivityLarge2024a, guoAutomatedPaperScreening2024b].
While prior studies show promising results for title-abstract screening in SE  [huotalaPromiseChallengesUsing2024a, felizardoChatGPTApplicationSystematic2024], they suffer from a small number of secondary studies and LLMs under evaluation. One study [huotalaPromiseChallengesUsing2024a] evaluated performance of two LLMs in a single secondary study, while another study [felizardoChatGPTApplicationSystematic2024] evaluated performance of a single LLM with two secondary studies. Performance evaluations based on small datasets risk introducing bias, making it difficult to assess the utility of the solution.
@@ II-A Automating the Screening Process of Systematic Reviews
[huotalaPromiseChallengesUsing2024a] SE 2 Binary 1 1,306
[felizardoChatGPTApplicationSystematic2024] SE 1 Likert 2 582
[liEvaluatingEffectivenessLarge2024] Biomed 9 Binary 3 505
[dennstadtTitleAbstractScreening2024] Biomed 3 Likert 10 38,426
[tranSensitivitySpecificityAvoidable2023] Med 1 Binary 5 22,666
[delgado-chavesTransformingLiteratureScreening2025] Med 18 Binary 3 6,217
[matsuiHumanComparableSensitivityLarge2024a] Med 2 Binary 2 4,527
[guoAutomatedPaperScreening2024b] Med 2 Binary 6 24,307
@@
Table I summarizes related work in this area. In the SE domain, two studies [huotalaPromiseChallengesUsing2024a, felizardoChatGPTApplicationSystematic2024] investigated the use of LLMs for screening with two different evaluation methods and 1,888 primary studies in total. In contrast, the biomedical and medical domains have seen more research on the topic [liEvaluatingEffectivenessLarge2024, dennstadtTitleAbstractScreening2024, tranSensitivitySpecificityAvoidable2023, delgado-chavesTransformingLiteratureScreening2025, matsuiHumanComparableSensitivityLarge2024a, guoAutomatedPaperScreening2024b]. Studies outside of SE have used significantly larger sample of both secondary and primary studies.
@@ II-B Benchmarking Large Language Models
AGIEval [liuAreLLMsGood2024, zhongAGIEvalHumanCentricBenchmark2023] Knowledge & reasoning
ARC-Challenge [clarkThinkYouHave2018] Knowledge & reasoning
BIG-Bench-Hard [suzgunChallengingBIGBenchTasks2023] Knowledge & reasoning
GPQA [reinGPQAGraduateLevelGoogleProof2023], Diamond Knowledge & reasoning
HellaSwag [zellersHellaSwagCanMachine2019] Knowledge & reasoning
MMLU [liuAreLLMsGood2024, hendrycksMeasuringMassiveMultitask2021] Knowledge & reasoning
GSM8K [liuAreLLMsGood2024, cobbeTrainingVerifiersSolve2021] Mathematical & logical reasoning
MATH [liuAreLLMsGood2024, hendrycksMeasuringMathematicalProblem2021] Mathematical & logical reasoning
MathVista (testmini) [luMathVistaEvaluatingMathematical2024] Mathematical & logical reasoning
MGSM [shiLanguageModelsAre2022] Mathematical & logical reasoning
MMMU [yueMMMUMassiveMultidiscipline2023] Mathematical & logical reasoning
HumanEval [liuAreLLMsGood2024, chenEvaluatingLargeLanguage2021] Programming & code generation
MBPP [austinProgramSynthesisLarge2021] Programming & code generation
WikiSQL [zhongSeq2SQLGeneratingStructured2017] Programming & code generation
DROP [duaDROPReadingComprehension2019] Multimodal & visual reasoning
AI2D, test [kembhaviDiagramWorthDozen2016] Multimodal & visual reasoning
ChartQA [masry-etal-2022-chartqa] Multimodal & visual reasoning
MMBench [liuMMBenchYourMultimodal2024] Multimodal & visual reasoning
@@ III Methods
Fig. 1 shows an overview of how the dataset was created and how the experiments used to answer our research questions were conducted. We followed the SIGSOFT Empirical Standards for Software Engineering Research [ralph2020empirical] as a framework for conducting this study. These standards provide guidance on reporting benchmarking studies.
@@ III-A Research questions
  • RQ4: What is the combined effect of LLMs and secondary studies on the screening results? Motivation: Understanding the combined effect of LLMs and secondary studies is important if certain LLMs perform significantly better in screening on specific secondary studies.
@@ III-B Creating the benchmark dataset
This subsection provides an answer for RQ1. Fig. 1 visualizes the benchmark dataset creation process. We used a prior systematic mapping [huotalaReplicationPackagesSoftware2025] as our source for research artifacts, which lists 169 research artifacts.
@@
From the data extraction (Table III), the ”Screening results” information is used to include or exclude research artifacts. We double-coded 20 randomized research artifacts regarding whether the secondary study contains screening data and that data is suitable for our experiment. We had an agreement of 90% and 95%, corresponding to Krippendorff’s alphas [krippendorff2011computing] of $0.786$ and $0.831$ respectively. The rest of the artifacts were then investigated by the first author.
@@
After the criteria retrieval, we generated unique identifiers for each secondary study in the dataset to distinguish them when running benchmarks. Finally, we combined the primary studies from each secondary study to form the final SESR-Eval dataset. Papers with missing titles or abstracts were excluded from the final dataset, as title-abstract screening is meant to be conducted with non-empty titles and abstracts. The final dataset is available in the research artifact of this study [zenodo_2_anon].
@@ III-C1 Creating the zero-shot prompt template and constructing the primary study prompts
Before constructing the primary study prompts, we created the zero-shot prompt template ( 8), in Fig.  2. As a starting point for the prompt template, we used the zero-shot prompt from a SE paper, which studied title-abstract screening using LLMs [huotalaPromiseChallengesUsing2024a]. Next, we added Likert scale decisions, used by Felizardo et al. [felizardoChatGPTApplicationSystematic2024], for each criterion (include or exclude) being met, and the relevance of the whole study:
@@
Finally, we added a floating point decision, which has a value between $0.000$ and $1.000$. A score of $1.000$ means the paper (or criterion) is included with high probability; secondly - a score of $0.000$ means the paper (or criterion) is most likely excluded. Our goal was to give the LLM as detailed prompt as possible, to reduce the chance of hallucinations and improve the overall task performance. We detailed each assessment type and emphasized that the LLM is required to produce all three types of assessments for the overall relevance of the study and for each individual inclusion-exclusion criteria.
@@ III-C2 Benchmarking title-abstract screening performance with LLMs
To study the screening performance of the LLMs with the larger dataset (RQ2) and across secondary studies (RQ3), we benchmarked the title-abstract screening performance with LLMs listed in Table IV. We aimed to include a diverse set of (a) reasoning, (b) non-reasoning, (c) commercial and (d) open-source LLMs, which publish their model weights or model source code online. For the reasoning models, we selected a commercial and an open-source LLM. Second, for non-reasoning models, we selected four commercial and two open-source LLMs. LLMs from these vendors are commonly used in SE research and are used in prior work [huotalaPromiseChallengesUsing2024a, felizardoChatGPTApplicationSystematic2024, houLargeLanguageModels2024, calciolariIntegratingLargeLanguage2025, delgado-chavesTransformingLiteratureScreening2025]. Similarly, as the performance of neural networks with a relatively small number of parameters, sometimes also called small language models (SMLs) can achieve competitive performance [schick2021s], we decided to add two models that could be considered such to better understand time and token efficiency related to model size. In all of our benchmark experiments, we used the following system prompt:
@@
With the LLMs, we used structured JavaScript Object Notation (JSON) response format instead of the traditional textual response. Using a structured response format allows for a predictable output token count, programmatic integration and faster evaluation of the LLM output [gengGeneratingStructuredOutputs2025, liuAreLLMsGood2024]. The response format we used is described in Table  V and available in the research artifact [zenodo_2_anon].
@@
To compare the performance of title-abstract screening between LLMs and secondary studies (RQ4), we fit a logistic regression model predicting the correctness of screening decisions based on the LLM and the primary study. The logistic regression model was fitted and the odds ratios were calculated using the R programming language and the methods glmglm and coefcoef included in the base R distribution version $4.3.3$. For calculating Krippendorff’s Alpha, we used the R-package irrirr [gamer2012package] version 0.84.1. We also calculated the precision, recall and the F1 score for the tested LLMs and all primary studies and secondary studies from the benchmark set, using Python’s scikit-learn library [pedregosa2011scikit]. The code for these is given in our research artifact.
For all LLMs, we set the temperature parameter  [^5] https://platform.openai.com/docs/api-reference/completions/create to zero and the top_p5 parameter to $0.1$ – configuration used in related works [huotalaPromiseChallengesUsing2024a, felizardoChatGPTApplicationSystematic2024]. The temperature parameter controls how deterministic the LLM’s output is. The lower the temperature value, the more the LLM attempts to always output the same tokens for the same prompt, regardless of the number of calls to the LLM. Secondly, the top_p parameter controls the top n-% tokens that are considered in the output. Setting top_p to $0.1$ means that the LLM attempts to sample from the top 10% most probable tokens, which guides the LLM to output more relevant tokens.
@@ III-D How does the time, token efficiency, and cost vary across different LLMs?
This subsection provides background for the last research question (RQ5). To evaluate the time, token efficiency and cost of the LLMs, we analyzed the latency and input and output token counts for each of the LLMs using a custom Python script. The latency and token information is returned by the LLMs’ API. Latency was measured as the duration of sending a single request to the LLM and receiving its response. We used the 95th percentile [hyndmanSampleQuantilesStatistical1996] of response times to evaluate performance under typical conditions. The total token count of the LLM is a combination of the input (=prompt) and output tokens, where reasoning models are expected to output additional ”reasoning tokens”, which are aggregated to the output tokens. Providers, such as OpenAI, cache input tokens to reduce costs  [^6] https://platform.openai.com/docs/guides/prompt-caching. Costs were computed from the token count, multiplied by the LLMs price per 1M tokens.
@@ IV Results
[stradowskiMachineLearningSoftware2023a] 1,194 742 451 62:38 0 Information and Software Technology 4 4 SE Process
[alonsoSystematicMappingStudy2023b] 223 191 32 86:14 1 Information and Software Technology 3 3 SE Process
[somersDigitaltwinbasedTestingCyber2023b] 458 146 312 32:68 53 Information and Software Technology 3 4 Software Testing
[sharbafConflictManagementTechniques2023a] 322 171 151 53:47 0 Software and Systems Modeling 3 6 SE Models and Methods
[zakeri-nasrabadiSystematicLiteratureReview2023b] 10,454 547 9,907 5:95 759 Journal of Systems and Software 4 6 Software Maintenance
[bjarnasonEmpiricallyBasedModel2023a] 4,671 54 4,617 1:99 968 Empirical Software Engineering 1 1 Software Design
[lewowskiHowFarAre2022a] 1,733 169 1,564 10:90 82 Information and Software Technology 13 0 Software Testing
[vandinterPredictiveMaintenanceUsing2022a] 606 144 462 24:76 47 Information and Software Technology 3 3 Software Maintenance / Operations
[casalaroModeldrivenEngineeringMobile2022a] 2,541 74 2,467 3:97 281 Software and Systems Modeling 6 4 Software Architecture
[deckersSystematicLiteratureReview2022a] 540 36 504 7:93 219 Journal of Systems and Software 3 6 Software Requirements
[tambonHowCertifyMachine2022a] 1,741 290 1,451 17:83 717 Automated Software Engineering 2 6 Software Quality
[raniDecadeCodeComment2023a] 2,353 73 2,280 3:97 274 Journal of Systems and Software 2 7 Software Quality
[pereiraLearningSoftwareConfiguration2021a] 69 69 0 100:0 0 Journal of Systems and Software 3 4 Software Configuration Management
[rodriguez-perezPerceivedDiversitySoftware2021a] 3,194 234 2,960 7:93 446 Empirical Software Engineering 7 7 SE Professional Practice
[barisicMultiparadigmModelingCyber2022a] 327 153 174 47:53 1 Journal of Systems and Software 5 6 SE Models and Methods
[ebrahimiMobileAppPrivacy2021a] 113 55 58 49:51 2 Information and Software Technology 3 3 Software Quality
[tebesAnalyzingDocumentingSystematic2020a] 731 522 209 72:28 236 Information and Software Technology 3 7 Software Testing
[geraldiSoftwareProductLine2020a] 167 56 111 34:66 0 Information and Software Technology 6 8 SE Models and Methods
[lenarduzziSystematicLiteratureReview2021a] 318 44 274 14:86 17 Journal of Systems and Software 4 7 Software Maintenance / SE Economics
[kuutilaTimePressureSoftware2020b] 5,454 223 5,231 4:96 1,268 Information and Software Technology 2 3 SE Management / SE Professional Practice
[shevtsovControlTheoreticalSoftwareAdaptation2018a] 1,512 161 1,351 11:89 25 IEEE Transactions on Software Engineering 4 0 SE Models & Methods
[teixeiraModelingAutomaticCode2017a] 318 187 131 59:41 0 Journal of Systems and Software 1 6 SE Models & Methods
[linOpinionMiningSoftware2022a] 802 127 675 16:84 21 ACM Trans. Softw. Eng 5 5 SE Professional Practice
[zhangTestingVerificationNeuralnetworkbased2020a] 105 105 0 100:0 1 Information and Software Technology 3 4 Software Testing
@@
Yes, 24 secondary studies were found with title-abstract screening data. The process of how the dataset was created is explained in Section III-B. In total, we reviewed 169 research artifacts, from which 11 (6.5%) were inaccessible due to reasons such as dead links, empty repositories or pages requiring authorization. Ultimately, 24 of the 169 artifacts (14.2%) were suitable for our benchmark dataset. The secondary studies selected for the benchmark dataset are shown in Table VI. The number of primary studies per secondary study varies significantly in the dataset — ranging from under a hundred papers to over 10,000 papers. Similar variations exist in the included/excluded ratio of papers. Two secondary studies contained only included studies, resulting in an I/E ratio of $100:0$, while at the other end of the spectrum, one study included only about one percent of papers, with an I/E ratio of $1:99$. On average, looking at the ratio of included and excluded studies in the dataset, 12% of primary studies were labeled as included and 88% as excluded. The median number of the inclusion and exclusion criteria across secondary studies was $3$ and $4.5$, respectively. Based on the Software Engineering Body of Knowledge (SWEBOK) [swebok], the dataset covers a wide spectrum of SE domains.